OpenStudy (anonymous):

simplify radical 6x^7y^11 The ^ sign are for exponents

OpenStudy (anonymous):

Does it mean? \[\sqrt{6x ^{7}y ^{11}}\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

\[\sqrt{6x ^{7}y ^{11}}=\sqrt{6x ^{6}x ^{1} y ^{10}y ^{1}}=\sqrt{6xy}*\sqrt{x ^{6}y ^{10}}\]

OpenStudy (anonymous):

\[x ^{3}y ^{5} \sqrt{6xy}\]